# Can Large Language Models Be Trusted as Black-Box Evolutionary Optimizers for Combinatorial Problems?

### v1

@@ 2.2 Error repair
The details of the different types of errors are listed in Table 1 (Please refer to the Appendix for more explanation) and an example of possible errors encountered in the population-level crossover is given in Figure 1(C). The outputs generated in different phases will undergo different validations due to the varying requirements for each phase. In addition, the output requirements are different when the input is an individual solution (denoted as $S$) and the entire population (denoted as $P$). The detail is shown in Table 2.
@@
In contrast, solutions with moderate errors that are not successfully repaired are allowed to enter into the next phase. This approach is adopted because (1) the error can always be mitigated after repair; and (2) moderate errors do not yield cascading effects on the entire optimization, and there will still be an opportunity to repair them in the next phases.
2.3 Solution evaluation
@@
3 Experimental studies
The fidelity and reliability of LLM-based EVO of different stages are validated on various datasets in different settings. The population size and the evolution round are set to 30. The number of tests for candidate selection is 100, and the results of evolutionary optimization are averaged from 10 independent simulations. The structural information of the tested dataset is listed in Table 3. More details of the parameters are given in the Appendix.
Dataset Dolphins Netscience Erods Email Astro
$\boldsymbol{|V|}$ 62 379 433 1,005 14,845
$\boldsymbol{|E|}$ 159 914 1,314 25,571 119,652
Table 3: The structural information of network dataset.
3.1 Optimization validation of LLM-based EVO
3.1.1 Initialization:
Table 4 shows the results of pass ratio of format and critical check in the candidate selection, consisting of ranking and filtering. As seen, LLMs effective on small datasets fail to generate valid outputs for large datasets. For example, on Dolphins, LLMs achieve 100.0% $\mathsf{Q}_{\mathsf{acc}}$ on both checks. We compare it with larger networks and observe an obvious decrease in pass ratios; particularly, the critical check pass drops to about 40%. Thus, it can be seen that LLMs’ performance in calculation tasks, like candidate selection, will decline as the amount of input data increases.
@@ 3.1.1 Initialization:
Table 4: The validation of LLM-based candidate selection.
The LLM-based initialization shows superior fitness at the starting point and during the optimization compared to random initialization, as shown in Figure 2. This result shows that LLMs can effectively handle population initialization, reducing human effort for networks of hundreds of nodes.
Figure 2: The optimization of different initialization methods.
3.1.2 Selection:
Figure 3 compares the selection performance of LLM-based selection with other selection strategies. The LLM-based selection demonstrates the highest fitness across the entire optimization generation, with only the tournament strategy achieving comparable performance. It can be deduced that LLMs is effective in decision-making tasks such as in the selection phase.
Figure 3: The optimization of different selection methods.
3.1.3 Reproduction:
@@
3.2 Reliability analysis
Netscience Format Error (GPT-4.0) Critical Error (GPT-4.0) Format Error (GPT-3.5) Critical Error (GPT-3.5)
@@ 3.2 Reliability analysis
Table 5: The validations of format and critical errors for the LLMs output in different phases.
Table 5 shows that the ratio of $\mathsf{Q}_{\mathsf{app}}$ regarding format error is consistently above 95% in all phases. Meeting the requirement of critical errors is more challenging as more criteria are involved. GPT-4.0 performs exceptionally well, achieving a $\mathsf{Q}_{\mathsf{rej}}$ below 1%. Even when issues are occasionally detected, they can be repaired, as seen in the initialization case. In contrast, LLM-based EVO enhanced by GPT-3.5 has a relatively low $\mathsf{Q}_{\mathsf{acc}}$ and has nearly 10% $\mathsf{Q}_{\mathsf{rej}}$ in critical error during the population-level crossover. Figure 5 also exhibits the similar result as format and critical check. The ratio of $\mathsf{Q}_{\mathsf{acc}}$ in moderate error of GPT-4.0 is higher than that obtained by GPT-3.5. It can be concluded that the quality of LLMs output is promising but strongly dependent on model capability. In phases that demand processing large data amount, such as population-level crossover and mutation, the LLM-based operator only achieves a maximum accuracy ($\mathsf{Q}_{\mathsf{acc}}$) of 92.3%. However, in phases requiring less data processing, it can maintain a minimum accuracy of 95%. There are also fewer errors for individual-level EVO than population-level EVO. We can conclude that the reliability of LLMs is mainly affected by the scale of the processed data.
Operation Dataset Format Error (GPT-4.0) Critical Error (GPT-4.0) Moderate Error (GPT-4.0)
@@
Table 6: The validations of format and critical errors for the LLMs output during crossover and mutation.
3.3 Scalability analysis
The scalability of LLM-based operators is subjected to the input data, which varies across different stages.
Initialization: As illustrated in Table 4, the effectiveness of candidate selection is significantly influenced by the size of the dataset. The input scale for LLMs during initialization (both candidate selection and sampling) increases linearly with the dataset size. For example, given a network of 10,000 nodes, we must input all their IDs with any specific metrics to LLMs for ranking, filtering, and sampling. It is not practical to input such large amounts of numerical data into LLMS and have them do such complex operations, not to mention the huge token overhead, meaning that LLMs may not be suitable for this data-intensive task.
@@ 3.3 Scalability analysis
Selection: No matter what the dataset is, the input to LLM is always $\{(S_{1},f_{1}),...,(S_{k},f_{k})\}$. Thus, the relevant factor to LLM-based selection is the population size. A larger population size increases the amount of data LLMs need to process, but the common setting of this parameter is manageable to LLMs [29, 24]. Therefore, LLM-based selection has great applicability in this decision-making task given the promising result in Figure 3.
Crossover and Mutation: For these two phases, the input to LLMs is either $[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}]$ (individual-level) or $\{[\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],\ldots,[\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}]\}$ (population-level). As such, the input data amount, the main factor affecting LLMs’ performance, is dependent on the solution size and population size (only applicable to population-level optimization). The population size and solution size are empirical and will unnecessarily increase with the increase in the scale of datasets. Thus, we can conclude that the dataset scale and the LLM’s performance are not directly related, which is demonstrated in Table 6 where the reliability of LLM-based EVO in the large dataset Astro is comparable to that in the two smaller networks.
Evidence of LLMs sensitivity to hyperparameters can be found in Table 7, where the population size is set to $\mathcal{P}_{1}=30$ and $\mathcal{P}_{2}=10$, and the solution size is set to $\mathcal{S}_{1}=10$ and $\mathcal{S}_{2}=5$. Reducing either the population size or the size of a single solution helps to minimize errors, a trend consistent for both crossover and mutation phases.
One potential factor affecting performance as dataset size increases is the representation of element IDs in the solution. For example, given two datasets with element counts of up to $10^{3}$ and $10^{4}$, the maximum number of digits required to represent element IDs is 4 and 5, respectively. It means that the growth in element ID digits increases logarithmically rather than proportionally with dataset size. Therefore, we can conclude the scalability of LLM-based EVO in the reproduction phase is excellent, which is also demonstrated in Table 6, where there is no obvious difference between Astro and other smaller datasets regarding three types of check.
Netscience Format Error (GPT-4.0) Critical Error (GPT-4.0) Moderate Error (GPT-4.0)
@@
Table 7: The validations for different errors for the LLMs output with different settings during crossover and mutation.
3.4 Ablation study
@@
3.5 Computational overhead analysis
@@
Figure 7: The computational cost of different cases.
4 Conclusion

### v2

@@ III-B3 Moderate Error
An example of possible errors encountered in the population-level crossover is given in Figure 1(C). The outputs generated in different phases will undergo different validations due to the varying requirements for each phase. In addition, the output requirements are different when the input is an individual solution (denoted as $S$) and the entire population (denoted as $P$). The detail is shown in Table II.
@@
In contrast, solutions with moderate errors that are not successfully repaired are allowed to enter into the next phase. This approach is adopted because (1) the error can always be mitigated after repair; and (2) moderate errors do not yield cascading effects on the entire optimization, and there will still be an opportunity to repair them in the next phases. For the process of validation and repair, please refer to Algorithm 3. In each stage, the output of LLMs will undergo the corresponding check and repair (if any). For the checklist of each stage, please refer to Table 2 of the main text.
Algorithm 3 CheckAndRepair
1:Output $\mathcal{O}_{x}$ obtained by $\mathbf{F}_{x}\in\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\}$
2:Refined output
3:Retrieve the checklist $\mathbf{L}_{F}$ in the phase regarding $\mathbf{F}_{x}$
4:for each $E_{x}$ in $\mathbf{L}_{F}$ do
5:     Check the input regarding $E_{x}$
6:     if $E_{x}$ passed then
7:         $\mathcal{O}_{x}\leftarrow\mathcal{O}_{x}$triangleright Output $\mathcal{O}_{x}$ remains unchanged
8:     else
9:         Retrieve the corresponding repair prompt $R_{x}$
10:         $\mathcal{O}_{x}\leftarrow\mathcal{F}_{R}(\mathcal{O}_{x},E_{x},R_{x})$triangleright pass the repaired result to the next check
11:     end if
12:end for
13:Return the final output $\mathcal{O}_{x}$
[…]
TABLE III: The prompt of population-level mutation
[…]
TABLE IV: The prompt of individual-level mutation
III-C Solution evaluation
@@
IV Experimental studies
In this section, we will examine the ability of LLMs to manipulate the solution of the network-structured problems.
IV-A Experimental settings and dataset
The fidelity and reliability of LLM-based EVO of different stages are validated on various datasets in different settings. The population size and the evolution round are set to 30. The number of tests for candidate selection is 100, and the results are averaged from 10 independent simulations.
We investigated using the population as the fundamental optimization unit, where all solutions undergo crossover and mutation. To fairly compare the population and individual-level optimization by LLM, we thus set the crossover and mutation rate to 1.0 for individual-level optimization. The ablation study employs GPT-4o, with initialization and selection done using GPT-4, while other tests utilize both GPT-4 and GPT-3.5. Other experimental settings are shown in Table V.
Parameter Value
LLM model GPT-3.5/4/4o
LLM temperature 0.8
Maximum token 1400
Population size 30
Number of generations per simulation 30
Solution size 10
Crossover rate 1.0
Mutation rate 1.0
Candidate number for initialization 50%
Node metric for candidate selection Degree
TABLE V: The experimental setting of LLM-based EVO.
Table VI provides more details about the tested networks, including $|V|$ and $|E|$ for the number of nodes and edges, respectively, $\langle K\rangle$ for the average degree, and CC and ASD for the clustering coefficient and average shortest distance. The data is available online [^1] http://www-personal.umich.edu/mejn/netdata..
Network $|V|$ $|E|$
Dolphins 62 159
Netscience 379 914
Erods 433 1,314
Email 1,005 25,571
Astro 14,845 119,652
TABLE VI: Topological information of networks.
IV-B Optimization problem formulation
The broad utility of evolutionary optimization for tackling discrete, non-linear problems has led to their extensive adoption in combinatorial challenges [73, 74]. In this study, we use the problem influence maximization that is extensively explored by the evolutionary computation community as an illustration [75, 57, 76] due to its generalizability. Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. Let $\{C_{1},C_{2},\ldots,C_{k}\}$ be the communities partition. For each community $C_{i}$, we can evaluate the influence of seed node set $S$ by calculating the intersection of their influenced nodes $I(S)$ with the nodes in $C_{i}$ as follows:
$I_{i}(S)=|I(S)\cap C_{i}|,$ (9)
where $I(S)$ refers to the function that simulates the spread of influence from the initial set $\mathcal{S}$.
The weight of each community $w_{i}$ is proportional to its size relative to the network. The fitness function $F(S)$ is then calculated as the weighted sum of the influences within each community, i.e.,
$f(S)=\sum_{i=1}^{k}\frac{|C_{i}|}{|V|}\cdot|I(S)\cap C_{i}|.$ (10)
IV-C Optimization validation of LLM-based EVO
IV-C1 Initialization
Table VII shows the results of the pass ratio of format and critical check in the candidate selection, consisting of ranking and filtering. As seen, LLMs effective on small datasets fail to generate valid outputs for large datasets. For example, on Dolphins, LLMs achieve 100.0% $\mathsf{Q}_{\mathsf{acc}}$ on both checks. We compare it with larger networks and observe an obvious decrease in pass ratios; particularly, the critical check pass drops to about 40%. Thus, it can be seen that LLMs’ performance in calculation tasks, like candidate selection, will decline as the amount of input data increases.
While LLMs possess the potential for global awareness, enabling them to generate a more diverse and well-distributed initial population, practical limitations arise when applied to large-scale networks. To make informed decisions, LLMs require structured input that includes node IDs along with relevant metrics, such as degree, betweenness centrality, or clustering coefficients. However, encoding this information for all nodes in a large graph can quickly exceed the model’s token limit, making full-graph initialization infeasible. This token overhead significantly restricts scalability, necessitating pre-filtering strategies or metric-based candidate selection before LLM engagement.
@@ IV-C1 Initialization
TABLE VII: The validation of LLM-based candidate selection.
Therefore, we only test the LLM-based initialization in the small networks as a preliminary study. The LLM-based initialization shows superior fitness at the starting point and during the optimization compared to random initialization, as shown in Figure 2. This result shows that LLMs can effectively handle population initialization, reducing human effort for networks of hundreds of nodes.
Figure 2: The comparison of different initialization methods.
IV-C2 Selection
The following are selection strategies for the comparison with LLM-based selection. Random selection chooses individuals randomly as parents for the next generation, without considering their fitness. No selection means every individual in the population survives to the next generation, regardless of fitness. Roulette selection assigns selection probabilities based on fitness, with fitter individuals having a higher chance of being selected. Tournament selection involves selecting a subset of individuals randomly and choosing the best from this subset as a parent. The process is repeated as needed.
Figure 3 compares the selection performance of LLM-based selection with other selection strategies. The LLM-based selection demonstrates the highest fitness across the entire optimization generation, with only the tournament strategy achieving comparable performance. It can be deduced that LLMs are effective in decision-making tasks such as in the selection phase.
Figure 3: The comparison of different selection methods.
IV-C3 Reproduction
@@
IV-D Reliability analysis
Netscience Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
@@ IV-D Reliability analysis
TABLE VIII: The validations for the output generated in different phases. The tested network is Netscience. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Erods Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
Initialization 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 92.0% 8.0% 0.0% 78.0% 14.0% 8.0%
Selection 100.0% 0.0% 0.0% 92.2% 5.5% 2.3% 100.0% 0.0% 0.0% 91.1% 6.9% 2.0%
Crossover (P) 99.3% 6.3% 0.1% 98.6% 1.4% 0.0% 93.3% 6.3% 0.4% 74.6% 13.4% 12.0%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.9% 0.1% 0.0% 100.0% 0.0% 1.0%
Mutation (P) 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 98.0% 1.6% 0.4% 93.0% 3.4% 3.6%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 98.0% 1.6% 0.4% 97.0% 1.3% 1.7%
TABLE IX: The validations for the output generated in different phases. The tested network is Erods. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Email Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
Initialization 100.0% 0.0% 0.0% 96.0% 4.0% 0.0% 98.0% 2.0% 0.0% 94.0% 6.0% 0.0%
Selection 100.0% 0.0% 0.0% 92.2% 5.5% 2.3% 100.0% 0.0% 0.0% 91.1% 6.9% 2.0%
Crossover (P) 98.0% 2.0% 0.0% 94.3% 4.3% 1.4% 96.3% 3.7% 0.0% 71.6% 10.4% 18.0%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.7% 0.2% 0.1% 100.0% 0.0% 1.0%
Mutation (P) 99.3% 0.6% 0.0% 99.7% 0.3% 0.0% 96.6% 3.4% 0.0% 89.8% 3.6% 6.6%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 94.4% 4.9% 0.7% 100.0% 0.0% 0.0%
TABLE X: The validations for the output generated in different phases. The tested network is Email. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Table VIII shows that the ratio of $\mathsf{Q}_{\mathsf{app}}$ regarding format error is consistently above 95% in all phases. Meeting the requirements of critical errors is more challenging as more criteria are involved. GPT-4 performs exceptionally well, achieving a $\mathsf{Q}_{\mathsf{rej}}$ below 1%. Even when issues are occasionally detected, they can be repaired, as seen in the initialization case. In contrast, LLM-based EVO enhanced by GPT-3.5 has a relatively low $\mathsf{Q}_{\mathsf{acc}}$ and has nearly 10% $\mathsf{Q}_{\mathsf{rej}}$ in critical error during the population-level crossover. Figure 5 also exhibits the similar result as format and critical check. The ratio of $\mathsf{Q}_{\mathsf{acc}}$ in moderate error of GPT-4 is higher than that obtained by GPT-3.5. It can be concluded that the quality of LLMs output is promising but strongly dependent on model capability. In phases that demand processing large data amount, such as population-level crossover and mutation, the LLM-based operator only achieves a maximum accuracy ($\mathsf{Q}_{\mathsf{acc}}$) of 92.3%. However, in phases requiring less data processing, it can maintain a minimum accuracy of 95%. There are also fewer errors for individual-level EVO than population-level EVO. We can conclude that the reliability of LLMs is mainly affected by the scale of the processed data.
Operation Dataset Format Error (GPT-4) Critical Error (GPT-4) Moderate Error (GPT-4)
@@
TABLE XI: The validations of format and critical errors for the LLMs output during crossover and mutation.
IV-E Scalability analysis
The scalability of LLM-based operators is influenced the input data, which varies across different stages.
IV-E1 Initialization
As illustrated in Table VII, the effectiveness of candidate selection is significantly influenced by the size of the dataset. The input scale for LLMs during initialization (both candidate selection and sampling) increases linearly with the dataset size. For example, given a network of 10,000 nodes, we must input all their IDs with any specific metrics to LLMs for ranking, filtering, and sampling. It is not practical to input such large amounts of numerical data into LLMS and have them do such complex operations, not to mention the huge token overhead, meaning that LLMs may not be suitable for this data-intensive task.
@@ IV-E1 Initialization
IV-E2 Selection
No matter what the dataset is, the input to LLM is always $\{(S_{1},f_{1}),...,(S_{k},f_{k})\}$. Thus, the relevant factor to LLM-based selection is the population size. A larger population size increases the amount of data LLMs need to process, but the common setting of this parameter is manageable to LLMs [77, 59]. Therefore, LLM-based selection has great applicability in this decision-making task given the promising result in Figure 3.
IV-E3 Crossover and Mutation
For these two phases, the input to LLMs is either $[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}]$ (individual-level) or $\{[\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],\ldots,[\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}]\}$ (population-level). As such, the input data amount, the main factor affecting LLMs’ performance, is dependent on the solution size and population size (only applicable to population-level optimization). The population size and solution size are empirical and will unnecessarily increase with the increase in the scale of datasets. Thus, we can conclude that the dataset scale and the LLM’s performance are not directly related, which is demonstrated in Table XI where the reliability of LLM-based EVO in the large dataset Astro is comparable to that in the two smaller networks. Evidence of LLMs sensitivity to hyperparameters can be found in Table XII, where the population size is set to $\mathcal{P}_{1}=30$ and $\mathcal{P}_{2}=10$, and the solution size is set to $\mathcal{S}_{1}=10$ and $\mathcal{S}_{2}=5$. Reducing either the population size or the size of a single solution helps to minimize errors, a trend consistent for both crossover and mutation phases.
One potential factor affecting performance as dataset size increases is the representation of element IDs in the solution. For example, given two datasets with element counts of up to $10^{3}$ and $10^{4}$, the maximum number of digits required to represent element IDs is 4 and 5, respectively. It means that the growth in element ID digits increases logarithmically rather than proportionally with dataset size. Therefore, we can conclude the scalability of LLM-based EVO in the reproduction phase is excellent, which is also demonstrated in Table XI, where there is no obvious difference between Astro and other smaller datasets regarding three types of check.
Netscience Format Error (GPT-4) Critical Error (GPT-4) Moderate Error (GPT-4)
@@ IV-E3 Crossover and Mutation
TABLE XII: The validations for different errors for the LLMs output with different settings during crossover and mutation.
IV-F Validation of moderate error
We have conducted an investigation about what specific moderate errors the LLM-base EVO would encounter during crossover and mutation. Only the population-level optimization is tested as individual-level optimization rarely produces errors, as observed from Table VIII and X. The errors can be found in Tables 1 and 2 in the main text. Here we list them out for reference.
  • Error 1: The population after the operation remains the same as before.
  • Error 2: The size of the population fails to meet the requirement.
  • Error 3: The new solution does not preserve the original number of nodes in the original solution.
  • Error 4: The new solution contains duplicated elements.
  • Error 5: The new solution contains invalid nodes not found in candidate nodes. (only applicable to mutation)
These errors will be examined sequentially according to their index. When error 1 occurs, it will request a new operation from LLM. For the rest of errors, we will attach the corresponding message to the newly generated population and input them to LLM for a repaired population. During the repair process, the population with moderate error from the last check will pass through to the next phase if format error or critical error occurs. Figure 6 presents the results of moderate error in mutation.
As observed, the selection of LLM is one of the main factors of the reliability. GPT-4’s improvement over GPT-3.5 in managing mutation errors is evident across all error types except error 4, with a marked increase in the proportion of solutions that meet quality standards without the need for further modifications. On the other hand, GPT-3.5 is more likely to produce the unacceptable population with format or critical errors especially when trying to repair error 1 and error 3 while it rarely happens for GPT-4. It can also be found the frequency of error varies, for example, error 3 and error 5 occur less frequently than others. Regardless of the LLM, error 1 happens most frequently although it can be repaired by GPT-4, it implies that LLM sometimes will be overwhelmed with the data and does not work at all. The increased effectiveness in handling errors by GPT-4 suggests advancements in the model’s capability to generate more robust and reliable solutions, possibly due to improved understanding and processing of complex scenarios.
Figure 7 illustrates the moderate errors encountered during the crossover of two different-sized networks, Netscience and Astro. In contrast to mutation (Figure 6), the $\mathsf{Q}_{\mathsf{acc}}$ ratio for LLM-based crossover is lower, suggesting that LLMs struggle more with crossover, possibly due to the need to manage two populations compared to the single population in mutation. We found that both networks exhibit a similar pattern despite their significant size difference, suggesting that network size has little impact on the performance of LLM-based EVO regarding reproduction and highlighting its excellent scalability.
Figure 6: The encountered moderate errors during population-level LLM-based mutation. The tested network is Netscience $(|V|=379)$.
Figure 7: The encountered moderate errors during population-level LLM-based crossover. GPT-4 is used as the LLM for testing. Two networks of different sizes are tested: Astro $(V=14,845)$ and Netscience $(|V|=379)$.
IV-G Ablation study of repair mechanism
@@
IV-H Computational overhead analysis
@@
Figure 9: The computational cost of different optimization modes.
V Conclusion